# Enhancing Reasoning Skills in Small Persian Medical Language Models Can Outperform Large-Scale Data Training

### v1

@@ I Introduction
In 2019, at the NeurIPS conference, Yoshua Bengio highlighted a critical limitation of current deep learning systems: their inability to perform tasks requiring robust reasoning skills [1]. While these systems excel at intuitive, perception-driven tasks, they struggle with reasoning-based challenges that demand deeper cognitive processing. This observation echoes Daniel Kahneman’s distinction between “fast” and “slow” thinking in cognitive science [2]. Fast, intuitive thinking aligns with the strengths of current AI systems, whereas slow, deliberate reasoning remains a significant weakness.
One proposed solution to this deficiency is the revival of symbolic AI methods through neurosymbolic approaches. However, Bengio cautioned against such methods due to scalability issues. Instead, he urged the development of new architectures capable of out-of-distribution generalization—a core requirement for genuine reasoning.
The Transformer architecture [3], which underpins most modern language models, suffers from the same deficiency. Transformers often make “stupid mistakes” when faced with reasoning-intensive problems, prompting researchers to enlarge models and datasets in an attempt to achieve better results across tasks. This approach arguably simulates System 2 reasoning within a fundamentally System 1 framework.
This limitation becomes even more pronounced in smaller language models, where restricted capacity and limited data exacerbate reasoning challenges. In domains such as medicine—where reasoning, interpretation, and decision-making are essential—these weaknesses become especially consequential.
Recent efforts (including ours) to enhance reasoning have largely focused on prompting or fine-tuning methods that enable models to imitate reasoning behavior rather than truly engage in it—a phenomenon we refer to as the “illusion of thinking” [4]. This means that such approaches provide only an approximation of genuine cognitive reasoning. Nevertheless, in the absence of architectures with intrinsic reasoning capabilities, these efforts represent necessary steps toward developing AI systems with more robust reasoning abilities.
The present work aims to address this challenge by enhancing the reasoning abilities of a small Persian language model, aya-expanse-8b [5]. We first prompt the model to generate Chain-of-Thought (CoT) trajectories for multiple-choice medical questions. These responses are then reviewed and corrected using DeepSeek-R, a model adept at reasoning, to rectify errors. This process constitutes Reinforcement Learning with AI Feedback (RLAIF) [6]. By pairing incorrect and corrected CoT trajectories, we subsequently apply Direct Preference Optimization (DPO) [7] to enhance the reasoning capabilities of the baseline model.
We call this improved model gaokerena-R. While its predecessor, gaokerena-V [8], demonstrated superior medical knowledge in straightforward prompting tasks, gaokerena-R outperformed it when evaluated using Chain-of-Thought (CoT) prompting [9]. Notably, gaokerena-R was trained on a much smaller dataset, focusing on reasoning enhancement rather than data scale, whereas gaokerena-V relied on extensive training on a large medical corpus and the MF3QA dataset [8].
Our findings highlight an important insight: in low-resource domains such as Persian medical AI, strengthening reasoning capabilities can be more effective than training on vast datasets. This underscores the importance of developing reasoning-oriented approaches, particularly in languages and fields with limited data availability. By improving reasoning abilities in smaller models, we can make meaningful progress toward practical, resource-efficient AI systems.
@@ II Related Work
To the best of our knowledge, no prior work has focused on developing Persian medical reasoning language models. Existing Persian medical models, including our previous gaokerena-V, mainly address knowledge representation and general language understanding with limited attention to reasoning.
L. Pan et al. reviewed several valuable studies on enhancing reasoning capabilities in language models. [10] Accordingly, we focus here on English-language works that offer methodological insights into generating and integrating reasoning data for improving medical language models.
@@ II-B Related Work In Other Domain
III Data
We required a Persian medical multiple-choice question answering dataset to evaluate our proposed methods. In the absence of such a dataset, we used DeepSeek-V3 [20] , a cost-effective large language model, to translate a small portion of the MedMCQA dataset from English to Persian. We randomly selected these questions from MedMCQA [21] to maintain topic diversity in the resulting dataset. To verify the quality of the translations, we prompted  [^1] Prompts are available at github.com/Mehrdadghassabi/Gaokerena-R/blob/main/dataset/judgement.ipynb two referees (grok-3-mini [22] and gpt-4.1-mini [23] ) to evaluate each translation. A translation was considered verified only if the both referees assigned it a score of 5 out of 5. This process resulted in approximately 11,000 Persian medical multiple-choice questions.
III-A PersianMedQA
Shortly after completing our translation process, a new Persian medical multiple-choice question answering dataset called PersianMedQA was introduced [24] . We randomly selected a small portion of its training data and appended it to our dataset to include additional data from a different source.
IV Proposed Methods
We proposed two methods. The first framework, depicted in 1, utilizes a teacher model ( we used DeepSeek-R), which is itself a reasoning language model, to correct the student model’s medical reasoning mistakes in its response to the given multiple choice question. (if the student model chose the correct option we move to the next question in our dataset) When provided with the correct answer, the teacher generates a detailed chain-of-thought explanation that highlights the reasoning behind the solution. In this setup, the teacher’s output serves as the preferred answer, while the student’s original response is treated as the rejected answer for Direct Policy Optimization (DPO).
The second framework, depicted in 2, employs the teacher model to provide more granular feedback by identifying specific errors in the student model’s answer in its response to the given multiple choice question. Through this constructive criticism, the student model is guided toward producing the correct answer independently. In this case, the final answer generated by the student is considered the preferred response, while its initial attempt functions as the rejected one for DPO. Although the second framework is considerably more time-consuming due to the iterative feedback process, it can significantly enhance the student model’s learning because DPO leverages preferred-rejected answer pairs generated solely by the student itself.
@@ IV Proposed Methods
Figure 1: Framework 1 Block Diagram
Figure 2: Framework 2 Block Diagram
V Results
@@
V-A Medical Reasoning Capabillities
To assess the medical reasoning capabilities of the models, we performed evaluations using the FA_MED_MMLU dataset  [^2] Available at huggingface.co/datasets/gaokerena/FA_MED_MMLU. Both gaokerena-V and gaokerena-R were prompted  [^3] Evaluation prompts are available at github.com/Mehrdadghassabi/Gaokerena-R/blob/main/evaluations/zeroshot-COT/kopp/gaokerena-r1.0/Untitled2.ipynb to generate Chain-of-Thought (CoT) reasoning traces with a temperature setting of 1.0 For each question, five independent samples were generated, and a majority-voting mechanism was applied to determine the final answer: if three or more of the five generations selected the same option, that option was chosen as the final prediction; otherwise, the question was left unanswered to reflect model uncertainty.
This evaluation framework provides a more robust estimation of reasoning consistency and agreement across multiple reasoning trajectories. The results are presented in Table I for the without negative marking setting, and in Table II for the with negative marking setting. These two scoring schemes allow for a fair comparison of the models’ reasoning accuracy under different evaluation criteria.
@@ V-A Medical Reasoning Capabillities
V-A1 Pass@K
The pass@k metric, originally introduced by B. Brown et al. [25], provides a robust measure for evaluating language models that generate varying outputs for the same input across multiple samples. It quantifies the probability that a model produces at least one correct answer within $k$ independent attempts, thereby offering a more comprehensive assessment of reasoning reliability and sample diversity. The formal definition of the metric is as follows, where $N$ denotes the total number of generated samples and $C_{i}$ represents the number of correct samples for problem $i$:
@@ V-A1 Pass@K
As illustrated in Figures 3 and 4, the gaokerena-R model consistently demonstrated superior performance across nearly all $k$ values and evaluation categories. This improvement indicates that gaokerena-R possesses more stable and coherent reasoning capabilities, producing correct answers more reliably even with a limited number of samples. In contrast, gaokerena-V exhibited relatively weak performance for smaller $k$ values, while its results improved as $k$ increased. This pattern suggests that gaokerena-V is considerably more uncertain when prompted to generate Chain-of-Thought reasoning trajectories, often producing diverse or inconsistent answers across different samples. The results therefore highlight the enhanced reasoning consistency and reliability achieved through gaokerena-R’s targeted reasoning-oriented training approach.
@@
V-B Medical Knowledge
As mentioned earlier, gaokerena-V possesses a broader base of medical knowledge compared to the newer gaokerena-R model, primarily because it was trained on a significantly larger and more diverse corpus of medical data. This difference becomes evident when both models are evaluated using direct question–answer prompting, rather than being asked to generate explicit Chain-of-Thought reasoning trajectories. Under this direct prompting setting, gaokerena-V demonstrates superior performance, reflecting its stronger memorization and factual recall abilities derived from large-scale medical pretraining. The comparative results for both models on the FA_MED_MMLU and IBMSEE (September 2023) datasets are presented in Table III, which clearly highlights gaokerena-V’s advantage in knowledge-intensive evaluations.

### v2

@@ I Introduction
Transformer-based language models [1] excel at fast, intuitive tasks—such as pattern matching, retrieval, and surface-level text generation—mirroring Kahneman’s concept of “fast thinking” [2]. However, they struggle with deliberate, multi-step reasoning tasks that require “slow thinking,” particularly in specialized domains such as medicine, where diagnostic accuracy depends on logical inference, evidence evaluation, and error correction. This limitation is even more severe in low-resource languages such as Persian, where both high-quality data and compute are scarce.
In 2019, Yoshua Bengio warned that deep learning systems lack true reasoning capacity and called for architectures that support out-of-distribution generalization [3]. While the transformer architecture was revolutionary, its success has largely come from scaling: larger models and larger datasets yield better performance. Yet, despite current large language models having trillions of parameters and being trained on tens of trillions of tokens, they still make surprisingly simple reasoning errors. They may even produce inconsistent answers when asked the same question directly or via chain-of-thought (CoT) prompting [4]. This deficiency becomes even more pronounced in small medical Persian language models, which have far fewer parameters and much less training data.
Recent advances have attempted to improve the reasoning ability of language models—enhancing their performance on “slow thinking” tasks—but these methods often rely on large, well-curated datasets available only for high-resource languages such as English or Chinese. In contrast, Persian lacks sufficient high-quality medical datasets. To address this gap, we propose a new framework to improve the reasoning capabilities of Persian medical language models under limited data availability.
In our proposed method, we machine-translated an English multiple-choice medical question answering dataset into Persian and applied reinforcement learning from AI feedback (RLAIF) [5] and direct preference optimization (DPO) [6] to enhance the reasoning ability of a baseline Persian medical model. The resulting model is named gaokerena-R.  [^1] All of our work is open-source and available at github.com/Mehrdadghassabi/Gaokerena-R
Our prior work, gaokerena-V [7], fine-tuned a Persian medical language model using approximately 57 million tokens (including a portion of a medical corpus and a dataset) via supervised fine-tuning (SFT). Although gaokerena-V demonstrated strong medical knowledge, gaokerena-R outperforms it when given chain-of-thought prompts—despite being trained on less medical data. Since both models share the same baseline model, aya-expanse-8b [8], we hypothesize that enhancing reasoning skills is more beneficial than scaling data for small, low-resource medical language models.
In summary, our contributions are as follows:
  1. 1.
    An efficient RLAIF+DPO framework that generates high-signal CoT preference pairs using a teacher–student loop.
  2. 2.
    gaokerena-R  [^2] Available at huggingface.co/gaokerena/gaokerena-r1.0 , an 8b-parameter medical model demonstrating that reasoning-focused training outperforms data scaling in low-resource medical NLP.
  3. 3.
    A machine-translated Persian medical multiple-choice question answering dataset designed for reasoning-focused model training.
@@ II Related Work
To the best of our knowledge, no prior work has focused on developing Persian medical reasoning language models. Existing Persian medical language models, including gaokerena-V, mainly address knowledge representation and general language understanding with limited attention to reasoning.
L. Pan et al. reviewed several valuable studies on enhancing reasoning capabilities in language models [9]. Accordingly, we focus here on English-language works that offer methodological insights into generating and integrating reasoning data for improving medical language models.
@@ II-B Related Work In Other Domain
III Proposed Methods
Two methods has been porposed. the first one, depicted in Figure 1, utilizes a teacher model (we used DeepSeek-R), which is itself a reasoning language model, to correct the student model’s medical reasoning mistakes in its response to the given multiple choice question. (if the student model chose the correct option we move to the next question in our dataset) When provided with the correct answer, the teacher generates a detailed chain-of-thought explanation that highlights the reasoning behind the solution. In this setup, the teacher’s output serves as the preferred answer, while the student’s original response is treated as the rejected answer for Direct Policy Optimization (DPO).
The second method, illustrated in Figure 2, employs the teacher model to provide more granular feedback by identifying specific errors in the student model’s answer to a given multiple-choice question. Upon receiving this feedback, the student model is prompted to correct its response precisely from the point where the error occurred. Through this iterative process, the student gradually refines its reasoning trajectory and converges on the correct answer. This constructive feedback allows the student model to develop reasoning independence, as it learns to correct its own mistakes rather than simply mimicking the teacher’s response.In this setup, the final answer generated by the student is considered the preferred response, while its initial attempt serves as the rejected one for DPO. Although this second framework is more time-consuming due to the iterative feedback process, it simplifies DPO training, since both the rejected and preferred answers are generated entirely by the student model itself as sequences of its own tokens.
@@ III Proposed Methods
Figure 1: Method 1 Block Diagram
Figure 2: Method 2 Block Diagram
IV Data
We required a Persian medical multiple-choice question answering dataset to implement our proposed method. In the absence of such a dataset at the time, we used DeepSeek-V3 [19], a cost-effective large language model, to translate a subset of the MedMCQA dataset from English to Persian. To maintain topic diversity, the questions were randomly selected from MedMCQA [20]. To verify the quality of the translations, we prompted  [^3] Prompts are available at github.com/Mehrdadghassabi/Gaokerena-R/blob/main/dataset/judgement.ipynb two referees—grok-3-mini [21] and gpt-4.1-mini [22]—to evaluate each translation. A translation was considered verified only if both referees assigned it a score of 5 out of 5. This process resulted in approximately 18,000 verified Persian medical multiple-choice questions.
IV-A PersianMedQA
Shortly after completing our translation process, a new Persian medical multiple-choice question answering dataset, called PersianMedQA, was introduced [23]. We randomly selected 1,000 questions from its training set and appended them to our dataset to incorporate additional data from a different source. The resulting dataset contained 19,000 Persian medical multiple-choice questions in total.
V Carbon Footprint
The carbon footprint of our DPO fine-tuning process was estimated based on the hardware configuration and total runtime. The procedure ran for a combined total of 1 hour on an NVIDIA H100 PCIe 80 GB GPU, with approximately 43 GB of VRAM utilized during training. The training loss curve is shown in Figure 3. Assuming an average power consumption of 350 watts per GPU, the total energy usage was approximately 0.35 kWh. Using the average carbon intensity of the Canadian electricity grid, where our server was located (0.086 kilograms of CO2 equivalent per kWh [24]), this corresponds to an estimated emission of 0.0301 kilograms of CO2 equivalent generated during the fine-tuning process. Compared to our previous model, gaokerena-V, which emitted 2.66 kilograms of CO2, this represents a substantial reduction in environmental impact.[25]
Figure 3: Training loss curve
VI Results
@@
VI-A Medical Reasoning Capabillities
To assess the medical reasoning capabilities of the models, we prompted them to produce Chain-of-Thought reasoning trajectories with a temperature setting of 1.0. Since all models can produce different answers for the same question  [^4] Evaluation prompts are available at github.com/Mehrdadghassabi/Gaokerena-R/blob/main/evaluations/zeroshot-COT/kopp/gaokerena-r1.0/Untitled2.ipynb, we evaluated them using two metrics on the FA_MED_MMLU  [^5] Available at huggingface.co/datasets/gaokerena/FA_MED_MMLU and IBMSEE (September 2023)  [^6] Available at huggingface.co/datasets/gaokerena/KOPP datasets.
VI-A1 Accuracy
The first metric is accuracy for each question in the datasets. Five independent samples were generated per question, and a majority-voting mechanism was applied: if three or more of the five generations selected the same option, that option was chosen as the final prediction; otherwise, the question was left unanswered to reflect model uncertainty. This framework provides a robust estimate of reasoning consistency across multiple trajectories. Results are presented in Table I (without negative marking) and Table II (with negative marking), where, in the negative marking setting, each incorrectly answered question is assigned a score of -0.33. These two scoring schemes allow for a fair comparison of the models’ reasoning accuracy under different evaluation criteria.
@@ VI-A1 Accuracy
VI-A2 Pass@K
The pass@k metric, originally introduced by B. Brown et al. [26], provides a robust measure for evaluating language models that generate varying outputs for the same input across multiple samples. It quantifies the probability that a model produces at least one correct answer within $k$ independent attempts, thereby offering a more comprehensive assessment of reasoning reliability and sample diversity. The formal definition of the metric is given in Formula 1 , where $N$ denotes the total number of generated samples and $C_{i}$ represents the number of correct samples for problem $i$.
@@ VI-A2 Pass@K
As illustrated in Figures 4 and 5, the gaokerena-R model consistently demonstrated superior performance across nearly all $k$ values and evaluation categories. This improvement indicates that gaokerena-R possesses more stable and coherent reasoning capabilities, producing correct answers more reliably even with a limited number of samples. In contrast, gaokerena-V exhibited relatively weak performance for smaller $k$ values, while its results improved as $k$ increased. This pattern suggests that gaokerena-V is considerably more uncertain when prompted to generate Chain-of-Thought reasoning trajectories, often producing diverse or inconsistent answers across different samples. The results therefore highlight the enhanced reasoning consistency and reliability achieved through gaokerena-R’s targeted reasoning-oriented training approach.
@@
VI-B Medical Knowledge
As mentioned earlier, gaokerena-V possesses a broader base of medical knowledge compared to the newer gaokerena-R model, primarily because it was trained on a significantly larger and more diverse corpus of medical data. This difference becomes evident when both models are evaluated using direct question–answer prompting, rather than being asked to generate explicit chain-of-thought reasoning trajectories. As shown in Table III, under this direct prompting setting, gaokerena-V demonstrates superior performance, reflecting its stronger memorization and factual recall abilities derived from large-scale medical pretraining. Interestingly, the approximately equal performance of gaokerena-R and its baseline aya-expanse-8b in this same direct prompting setup indicates that gaokerena-R’s superior results in the chain-of-thought setting stem primarily from its enhanced reasoning skills rather than from increased medical knowledge.
@@ VI-C Final Evaluation
VII Future Research
The results show that gaokerena-V performs better with direct prompts, while gaokerena-R excels with chain-of-thought (CoT) prompts. This indicates that both models are prompt-dependent, as their performance varies with the prompting style. Future research should aim to develop prompt-invariant medical language models that integrate strong reasoning skills and medical knowledge, achieving superior performance regardless of the prompt format. Achieving prompt invariance would represent an important step toward more reliable and generalizable small Persian medical language models.